Solve the system by the elimination method.
5x+3y=-11
7x-2y=13
Thank you to whoever can help me.

Answer by mollukutti(30)

Basic Facts:
For elimination method, coefficients of one of the variables i.e, x or y needs to be equalized first
In this case we will try to equalize the co-efficients of x

5x+3y=-11 ..... multiply the expressions on either sides by 7
7x-2y=13 ..... multiply the expressions on either sides by 5
______________
35x + 21y = -77........ now let us subtract the left hand side expression and the
35x - 10y = 65......... right hand side expression
_______________
(35x + 21y)-(35x - 10y) = (-77)-65
or, 35x + 21y - 35x + 10y = -77 -65
or, 35x - 35x + 21y +10y = -142
or, 31y = -142
or, y = -142/31
By substitution we find the value of x now
7x - 2y = 13
or, 7x - 2. -142/31 = 13
or, 7x + 284/31 = 13
or, 7x = 13 - 284/31
or 7x = 119/31
or, x = 17/31 (dividing either sides by 7)